OpenStudy (anonymous):

An aeroplane travels at an average speed of 600kmh-1 on an outward flight and at 400kmh-1 on the return flight over the same distance What is the average speed of the whole flight? A) 111ms-1 B) 167ms-1 C)480kmh-1 D)500kmh-1

OpenStudy (anonymous):

I know a blog for physics mcq"s you're doing alevel?

OpenStudy (anonymous):

Some may give the average as (600 + 400) / 2 = 500 km h-1. [Choice D] But this is wrong. Averages are not that simple unless the 2 times are the same and not the 2 distances. Average speed = Total distance / Total time Let distance travelled for outward flight, which is equal to the distance travelled for inward flight = d. So, total distance travelled for the outward and return flights = 2d. For each flight speed, v = d / (time, t). The time will be different since the speed is different for the flights. let time for outward flight = tout. tout = (d/600) h. [h: hour] let time for inward flight = tin. tin = (d/400) h. Total time, t = tout + tin = d/600 + d/400 = (d/240) h Average speed = total distance / total time = 2d / (d/240) = 480 km h-1

OpenStudy (anonymous):

^ this solution is copied its from http://physics-ref.blogspot.com a really great blog for physics alevel.

OpenStudy (anonymous):

Oh thankyou so much :') And yes i am doing alevels

OpenStudy (anonymous):

No problem.